# Edge Graph Intelligence: Reciprocally Empowering Edge Networks with Graph Intelligence

### v1

@@ III-C1 Software Frameworks for Edge Computing
Software frameworks for edge computing across edge networks are often cross-platform, cross-protocol, language-agnostic, and resource-efficient. EdgeX Foundry [94] is an open-source edge platform that lets users create IoT gateway functionality from edge devices, which acts as a dual transformation engine sending and receiving data to and from cloud and edge applications. Eclipse Kura [95] has taken a similar approach to EdgeX Foundry to provide a platform for developing IoT gateways. Apache Edgent [96] is an open-source stream processing framework for edge computing, enabling developers to process sensor-collected streaming data in real time on edge devices. RedisEdge [97] is a purpose-built, multi-model database for the demanding conditions at the IoT edge, which can ingest millions of writes per second with less than 1ms latency and a very small footprint (less than 5MB). TensorFlow Lite [98] and MNN [99] are an open-source lightweight framework that enables on-device machine learning by helping developers run their AI models on embedded and mobile devices to achieve edge intelligence.
III-C2 Virtualizing and Resource Management Techniques
The requirements of Virtualizing and resource management technology reflect in the following aspects: (1) The challenge of uniformly managing resources across widely distributed and diverse devices in edge networks. (2) The need for resource isolation on edge nodes running multiple applications to prevent contention. (3) Edge applications’ reliance on various software library versions and dependencies, where virtualization aids in efficient management and compatibility. KubeEdge [100] and Baetyl [101] are open-source container-centric virtualizing and resource management techniques based on Kubernetes [102]. They deploy applications on edge nodes in the form of containers and offer a cloud-based management suite to oversee all resources within the edge network. Azure IoT Edge [103] enables developers to package services into modules and run them using the Azure IoT Edge runtime. Multiple modules can be configured to communicate with each other, forming a data processing pipeline. AWS IoT Greengrass [104] enables the deployment, execution, and management of Docker containers on edge devices and facilitates connectivity to AWS services, bringing cloud processing and logic to edge devices.
III-C3 Simulation Toolkits
@@ III-D Edge Intelligence
In examining the advancements in the field of AI, particularly in the realms of DL model training and inference, it becomes clear that edge intelligence holds distinct advantages over traditional cloud intelligence: (1) Model Training: In the current AI landscape, there has been a significant shift in data sources from centralized cloud data centers to the increasingly ubiquitous edge devices, such as mobile and IoT devices. These methods require the collection of massive amounts of data in the cloud, exerting immense pressure on core networks due to high data traffic. Additionally, this centralized approach strains data center resources, both in terms of storage and data processing capabilities. Edge intelligence offers a strategic solution to this challenge. By facilitating the processing and training of deep learning models directly at the network edge, closer to where the data is generated [109, 110, 111]. (2) Model Inference: Upon completion of model training, these models are deployed to provide inference services to end users. Routing user requests to cloud-based systems for inference poses significant privacy concerns, as it could lead to the leakage of sensitive information, such as personal identifiers and location data. Moreover, relying on cloud-based inference services, which depends on the stability of core networks, may introduce significant propagation delays. This aspect is particularly critical for applications requiring real-time response, such as autonomous driving and medical-surgical robots. Edge intelligence, by processing these inference requests locally, not only enhances data privacy but also ensures low-latency, efficient service delivery, crucial for such time-sensitive applications [112, 113, 114]. EGI can be regarded as a particular genre of edge intelligence concerning GI and is yet in its seed time for exploration and exploitation. In the following sections, we will discuss EGI in detail from different aspects.
IV Graph Intelligent Applications at Edge
@@
Graphs are highly abstract data structures and can be used to represent a spectrum of data generated in edge networks. These graph data drive miscellaneous applications at the last mile of the Internet, as exemplified in Fig. 7, and catalyze graph learning as a promising principle for edge intelligence. In this section, we present representative GI-enabled edge applications and highlight their superior endowment for high-performance services. For clarity, we categorize them into five groups, i.e., smart cities, robots and vehicles, human sensing and analysis, location-based recommendation, and graph-based mobile vision.
IV-A Smart Cities
Managing large cities is one of the key problems for modernization, especially under the stresses of continuously booming urbanization and the rapidly growing population. To enable resource-efficient management, AI has been widely recognized as an advantageous way for assist, promoting the concept of “smart cities”. Given the ubiquity of wireless sensor networks and graph data in cities, GI shines among AI techniques and enables numerous smart applications. Note that many use cases may not be fully deployed on edge networks but involve the cloud as a coordinator, our review focuses on the processing of graph data collected at the edge and demonstrates how EGI works for them. In the following, we discuss representatives concerning transportation, energy, environment, and public safety.
IV-A1 Intelligent Transportation Systems
@@
Traffic state forecasting involves forecasting traffic amount, traffic flow, and travel time between two locations, etc. Typically, it takes data from roadside sensors and organizes them in a graph along with the structure of the road net, as visualized in Fig. 7(a), where the feature vector of each vertex is the collected data of each sensor. GI models can thereafter be utilized to learn the semantics of traffic networks and export predictions. For example, Li et al. [19] introduce the Diffusion Convolutional Recurrent neural network (DCRNN), an advanced deep learning framework designed to tackle the complexities of traffic forecasting. The DCRNN effectively models traffic flow as a diffusion process on a directed graph, capturing both spatial dependencies through bidirectional random walks and temporal dynamics using an encoder-decoder architecture. Also recognizing the critical role of spatial and temporal dependencies in traffic flow, Yu et al. [117] present Spatio-Temporal Graph Convolutional Networks (STGCN), a novel approach for time series prediction in traffic analysis. Instead of applying regular convolutional and recurrent units, STGCN utilizes a fully convolutional structure on graphs, enhancing training efficiency and reducing parameter count.
@@
Supply and demand predictions are a type of trip prediction task though they are often viewed as an a ensemble, namely the origin-destination demand prediction. This can be relevant in many ride-hailing services (e.g., Uber [118], DiDi [119], and Lyft [120]), where accurately predicting users’ demand is crucial for dispatch orders concerning efficiency and profit. Wang et al. [121] propose a novel spatio-temporal graph attention network, named Gallat, to capture the dynamic, directed, and weighted nature of passenger demand graphs. By incorporating three attention layers, Gallat effectively models the spatio-temporal dependencies among regions across historical time slots. Building on the importance of capturing dynamic dependencies in traffic flow, Lu et al. [122] developed a dual attentive GNN that can effectively predict the distribution of metro passenger flow considering the spatial and temporal influences. Further expanding on the application of attention mechanisms in GNNs, Makhdomi et al. [123] introduces a GNN framework incorporating the Attention Mechanism for passenger origin-destination flow prediction. The framework leverages diverse linear and non-linear dependencies among requests originating from distinct locations, capturing both the repetition pattern and contextual data of each location. Their superior performance demonstrates the capability to capture inherent connections between regions or origin-destination pairs, and some of them have been deployed in the wild.
@@
Traffic resource management is to detect or schedule dedicated components in transportation systems. As an example, Wang et al. [124] and Yu et al. [125] both introduce GNN-based learning frameworks aimed at addressing critical issues in intelligent transportation systems. These frameworks capture complex dependencies within their respective applications, showcasing the versatility of GNNs in this field. However, their focuses differ significantly. Wang et al. [124] propose an improved contrastive GNN-based learning framework specifically designed for traffic anomaly analysis in IoT-based intelligent transportation systems, aiming to identify and analyze traffic anomalies, thereby enhancing the overall safety and efficiency of transportation networks. In contrast, Yu et al. [125] focus on the operational problem of relocating and repositioning idle vehicles, a crucial aspect of traffic resource management.
IV-A2 Power Grid
@@
Failure detection aims at localizing potential or yet-occurring failures within the power grid. Liao et al. [126] introduce a transformer fault diagnosis method using a GCN, which combines the strengths of existing approaches by utilizing the GCN’s adjacency matrix to represent the similarity between unknown and labeled samples, and its convolutional layers for identifying complex nonlinear relationships between dissolved gas and fault type. Introducing GCN into power grids enhances diagnostic accuracy by integrating feature extraction with relational learning, outperforming conventional techniques that may overlook the underlying graph structure of power systems. Building on the use of GCNs in power system diagnostics, Liu et al. [127] further present a technique that employs GCNs for swiftly identifying key failure points in power systems. This method simplifies and speeds up the process of detecting critical cascading failures, making it highly effective for complex power networks. Extending the application of GCNs to the realm of cybersecurity in power grids, Boyaci et al. [128] introduce a deep learning model utilizing Chebyshev Graph Convolutional Networks (CGCN) for detecting cyberattacks, specifically false data injection attacks, in large-scale AC power grids. The CGCN model stands out by efficiently modeling the inherent graph structure of power grid measurements, leveraging spatial correlations to enhance detection accuracy.
Energy harvesting prediction is for power grids with vulnerable energy input, e.g., solar energy and wind energy. Accurately predicting their future behavior is of great importance for power systems with high penetration of RESs for stable operation and economic consideration. Karimi et al. [129] focuses on improving photovoltaic power forecasting by leveraging spatial and temporal coherence among power plants. The authors propose a spatio-temporal GNN to harness the relationship between power plants, observing that plants in a region experience similar environmental conditions and can thus inform each other’s power forecasts. Extending the use of GNN for renewable energy forecasting, Khodayar et al. [130] introduce a convolutional graph auto-encoder (CGAE) for capturing continuous probability densities on graph nodes. Employing spectral graph convolutions, deep learning, and variational Bayesian inference, CGAE generates samples from these densities. This approach is applied to probabilistic solar irradiance prediction, using an undirected graph model of solar radiation measurement sites for future irradiance estimation. In their subsequent research, Khodayar et al. [131] further tackle the challenge of forecasting behind-the-meter load and rooftop photovoltaic generation in power systems. It uses a spatio-temporal graph autoencoder and graph dictionary learning optimization to disaggregate historical net load data and predict future behind-the-meter load and photovoltaic generation.
IV-A3 Environment Monitoring
@@
Environmental assessment and prediction aims at extracting environmental information from sensory data collected across smart cities. Chen et al. [132] introduce the group-aware GNN (GAGNN) for nationwide air quality forecasting, in order to understand latent dependencies among geographically distant cities. GAGNN constructs a hierarchical model with a city graph and city group graph, combined with a differentiable grouping network to discover and encode these complex inter-city relationships. In addition to the use of GCNs, both Gao et al. [133] and Mao et al. [134] incorporate Long Short-Term Memory (LSTM) networks to enhance air quality prediction. By leveraging GCNs to model spatial dependencies and LSTM networks to capture temporal dynamics, they effectively integrate spatio-temporal information for more accurate predictions. Specifically, Gao et al. [133] introduce the graph-based long short-term memory (GISTM) model for predicting PM2.5. This model integrates an adjacency matrix within the LSTM cell, treating air quality monitoring stations as a graph to incorporate spatio-temporal information. Building upon this approach, Mao et al. [134] develop a graph convolutional temporal sliding LSTM model to predict various air pollutants. They use GCNs to model spatial dependencies and a temporal sliding LSTM strategy to capture dynamic changes over time, allowing for a broader application beyond just PM2.5 prediction.
Meteorological monitoring and forecasting leverage GI models to analyze meteorological data for understanding current weather conditions and predicting future weather patterns. The targeted sensory data are usually in a spatio-temporal form, covering atmospheric parameters such as temperature, humidity, pressure, wind speed and direction, precipitation, and other relevant variables. GCNs excel in understanding current weather conditions and predicting future patterns due to their ability to handle irregular spatial distributions of meteorological data. Unlike traditional grid-based models, GCNs represent the complex topology of the Earth’s surface, capturing interactions between geographically distant but climatically connected regions. Building on this strength, Lin et al. [135] present the Conditional Local Convolution Recurrent Network for spatio-temporal forecasting in weather prediction, overcoming challenges of high nonlinearity and complex spatial patterns. It introduces a GCN that captures local spatial patterns through conditional local convolutions and incorporates these into an RNN architecture to model temporal dynamics. Addressing the challenge of accurately forecasting frost events, which significantly impact agriculture, Lira et al. [136] describe a GNN with spatio-temporal attention architecture. The model, adept at handling the localized nature of frost influenced by various environmental factors, maps weather station data onto a graph structure, optimizing an adjacency matrix during training to capture complex environmental interactions. Similarly focusing on the spatial relationships among weather stations, Stanczyk et al. [137] introduce novel models for wind speed prediction utilizing GCNs, focusing on handling data from multiple weather stations. These stations are represented as nodes in a graph, with an adjacency matrix that is learnable, allowing the network to understand the spatial structure and interrelations based on historical weather data.
IV-A4 Public Safety

### v2

@@ III-B3 Software Frameworks for Edge Networks
Software frameworks for edge computing across edge networks are often cross-platform, cross-protocol, language-agnostic, and resource-efficient. EdgeX Foundry [78] is an open-source edge platform that lets users create IoT gateway functionality from edge devices, which acts as a dual transformation engine sending and receiving data to and from cloud and edge applications. Eclipse Kura [79] has taken a similar approach to EdgeX Foundry to provide a platform for developing IoT gateways. Apache Edgent [80] is an open-source stream processing framework for edge computing, enabling developers to process sensor-collected streaming data in real time on edge devices. RedisEdge [81] is a purpose-built, multi-model database for the demanding conditions at the IoT edge, which can ingest millions of writes per second with less than 1ms latency and a very small footprint (less than 5MB). TensorFlow Lite [82] and MNN [83] is an open-source lightweight framework that enables on-device machine learning by helping developers run their AI models on embedded and mobile devices to achieve edge intelligence.
@@ III-B4 Edge Intelligence
In examining the advancements in the field of AI, particularly in the realms of DL model training and inference, it becomes clear that edge intelligence holds distinct advantages over traditional cloud intelligence: (1) Model Training: In the current AI landscape, there has been a significant shift in data sources from centralized cloud data centers to increasingly ubiquitous edge devices, such as mobile and IoT devices. These methods require the collection of massive amounts of data in the cloud, exerting immense pressure on core networks due to high data traffic. Additionally, this centralized approach strains data center resources, both in terms of storage and data processing capabilities. Edge intelligence offers a strategic solution to this challenge. By facilitating the processing and training of deep learning models directly at the network edge, closer to where the data is generated [88, 89, 90]. (2) Model Inference: Upon completion of model training, these models are deployed to provide inference services to end users. Routing user requests to cloud-based systems for inference poses significant privacy concerns, as it could lead to the leakage of sensitive information, such as personal identifiers and location data. Moreover, relying on cloud-based inference services, which depend on the stability of core networks, may introduce significant propagation delays. This aspect is particularly critical for applications requiring real-time response, such as autonomous driving and medical-surgical robots. Edge intelligence, by processing these inference requests locally, not only enhances data privacy but also ensures low-latency, efficient service delivery, crucial for such time-sensitive applications [91, 92]. EGI can be regarded as a particular genre of edge intelligence concerning GI and is yet in its seed time for exploration and exploitation. In the following sections, we will discuss EGI in detail from different aspects.
IV Edge Graph Intelligence as a New Edge AI Paradigm
Stemed from edge AI, EGI dives into the fusion of edge networks and GI techniques, serving as a brand-new avenue for the evolution of AI deployment. In what follows, we discuss the benefits of EGI and provide a rating taxonomy of EGI in six levels.
IV-A Reciprocal Benefits of Edge Graph Intelligence
EGI provides reciprocal benefits in the following aspects.
  • Edge networks for GI. With the rapid proliferation of mobile and IoT devices, data generated at edge networks have skyrocketed in both quantity and modality (e.g., physical signals, digital audio, and visual content). As predicted by IDC [93], the billions of IoT devices in edge networks are expected to generate over 90 Zettabytes of data in 2025. This naturally provides rich data nurseries and abundant application scenarios for modifying, training, and fine-tuning GI models with real-world data, thus boosting GI models toward higher-degree intelligence. Besides, edge networks as user-nearby infrastructures can effectively enhance GI model computing performance, e.g., edge networks assisted with cloud as the back-end can largely reduce the computing latency, thus boosting the GI performance.
  • GI for edge networks. Given the rich relational data collected at edge networks, GI enables modern graph analysis for understanding, diagnosing, and optimizing edge networks, which steers the enhancement of network performance such as robustness and Quality of Service (QoS). Applying GI to edge networks thus unlocks their extended capability in securing edge networks from anomalies, developing new graph-based applications, and intelligently serving graph-related tasks.
  • Reciprocal technology advancement. The combination of GI and edge computing not only carries out mutual empowerment to improve each other (as mentioned above) but also enhances the advancement of both research and technology. For instance, EGI catalyzes a number of new edge applications (Sec. V) and incubates an ecosystem including GNN-oriented hardware accelerators and programming frameworks (Sec. VIII), which actually expedites edge AI popularization. Additionally, and perhaps more notably, new technologies and system models for edge AI are being nourished and flourishing, extending beyond the traditional scopes and standards of these two.
IV-B Rating of Edge Graph Intelligence
Given the reciprocal benefits of EGI, we advocate that EGI should not be restricted to merely applying GI on edge data or running GI on edge platforms. Instead, GI and edge networks are blending a confluence and EGI should be treated as a whole to reflect the inherent interplay between GI and edge networks. This indicates that their bilateral empowerment desires a comprehensive exploration such that the degree of EGI can be identified and measured. Specifically, according to the fusion of GI and edge networks, we may rate EGI into six levels from their respective perspectives, as shown in Fig. 6.
  • Level 0: Given the graph data implicated in the edge network infrastructure, analytical models in Level 0 are unaware of the graph structures. The edge computing systems also process data in a graph-agnostic way. In other words, neither the model side nor the infrastructure side explicitly tackles “graphs”, and thus they are categorized to the initial level. For instance, visual data mining with traditional ML models and conventional neural networks like CNN [94, 95] and RNN [96, 97] can be categorized in Level 0 since their data are not formulated in graphs and their data processing pipeline has no specific relations with graphs.
  • Level 1: Data collected from edge networks are modeled in graphs but are not exploited in analytics. Systems at Level 1 push one step further over Level 0 by endowing graph semantic to edge data through graph modeling with collected edge data, but process data with general computing methods (instead of graph-related methods). As an example, while wireless sensory data can be generally abstracted in graphs [37, 38], Level-1 systems treat them in a vertex-wise manner, e.g., separately analyzing individual sensors’ data arrivals, instead of processing as a whole graph.
  • Level 2: Edge data in graph formats are processed with traditional graph computing algorithms such as PageRank [98] and single-source shortest path algorithms [99]. In this respect, edge data are modeled in graphs and processed with traditional graph algorithms rather than ML methods. Systems at Level 2 outperform Level 1 by enabling graph-oriented computing capability though they do not introduce AI techniques in graph processing.
  • Level 3: Edge networks serve GI model inference with graph data, where the models may be trained on the cloud. Compared with lower levels, systems at Level 3 not only allow graph formulation and computing of edge data but also initiate AI to edge networks and embrace GI models such as GCN [15] and GraphSAGE [48]. In other words, edge systems that compute GNN inference such as GNN-based traffic forecasting systems [18, 100] and DGRL decision systems [101, 102] are located at Level 3.
  • Level 4: Edge networks perform GI model training with graph data. Example includes SUGAR [103] and HGNAS [104]. The key difference between Level 4 and Level 3 lies in the ability to learn edge-native GI models, e.g., fine-tuning model parameters with edge data. Otherwise stated, systems at Level 4 can customize their GI models in-situ, thereby enabling tailored ability for edge AI services.
The EGI rating can be mainly divided into three intervals. The first interval covers from Level 0 to Level 2, where EGI is less related to AI and even processes non-graph data. The second interval comprises Level 3 and Level 4, where EGI incorporates GI models by either inference or training on edge networks. The third interval is exactly Level 5, which stands at the highest level because its GI and edge networks have profoundly blended as integration and can adapt to diverse scenarios on the fly. As EGI systems are located at higher levels, their fusions of GI and edge networks go deeper. As a result, the intelligence resources of GI and infrastructural resources of edge networks are progressively exploited for better EGI performance. Nonetheless, this may also come at the cost of additional development effort and system overhead. This conflict implies that there is no “silver bullet” in all cases. Instead, the panacea of EGI in practice should align with user demand, anticipating a joint consideration of specific application scenarios as well as available resources budgets. In this survey, given the focus on advanced GNN in edge environments, the reviewed systems mostly lie in Level 3 and beyond.
V Edge Applications of Graph Intelligence
@@
Graphs are highly abstract data structures and can be used to represent a spectrum of data generated in edge networks. These graph data drive miscellaneous applications at the last mile of the Internet, as exemplified in Fig. 7, and catalyze graph learning as a promising principle for edge intelligence. Following the insights discussed in Sec. IV, we first provide a panorama of representative EGI applications before showing how edge networks and GI interact with each other in subsequent sections. For clarity, we categorize them into five groups, i.e., smart cities, robots and vehicles, human sensing and analysis, location-based recommendation, and graph-based mobile vision. Note that edge applications of GI models mostly only compute GI model inference at the edge and these related EGI systems therefore lie at Level 3 in the rating taxonomy.
V-A Smart Cities
Managing large cities is one of the key problems for modernization, especially under the stresses of continuously booming urbanization and the rapidly growing population. To enable resource-efficient management, AI has been widely recognized as an advantageous way to assist, promoting the concept of “smart cities” [107, 108]. Given the ubiquity of wireless sensor networks and graph data in cities, GI shines among AI techniques and enables numerous smart applications. Note that many use cases may not be fully deployed on edge networks but involve the cloud as a backend, our review focuses on the processing of graph data collected at the edge and demonstrates how these applications run. In the following, we discuss representatives concerning transportation, energy, environment, and public safety.
V-A1 Intelligent Transportation Systems
@@
Traffic state forecasting involves forecasting traffic amount, traffic flow, and travel time between two locations, etc. Typically, it takes data from roadside sensors and organizes them in a graph along with the structure of the road net, as visualized in Fig. 7(a), where the feature vector of each vertex is the collected data of each sensor. GI models can thereafter be utilized to learn the semantics of traffic networks and export predictions. For example, Li et al. [18] introduce the Diffusion Convolutional Recurrent neural network to model traffic flow as a diffusion process on a directed graph, capturing both spatial dependencies through bidirectional random walks and temporal dynamics using an encoder-decoder architecture. Recognizing the critical role of spatial and temporal dependencies in traffic flow, Yu et al. [100] present Spatio-Temporal Graph Convolutional Networks, a novel approach for time series prediction in traffic analysis. Instead of applying regular convolutional and recurrent units, STGCN utilizes a fully convolutional structure on graphs, enhancing training efficiency and reducing parameter count.
@@
Supply and demand predictions are a type of trip prediction task though they are often viewed as an ensemble, namely the origin-destination demand prediction. This can be relevant in many ride-hailing services (e.g., Uber [109], DiDi [110], and Lyft [111]), where accurately predicting users’ demand is crucial for dispatch orders concerning efficiency and profit. Building on the importance of capturing dynamic dependencies in traffic flow, Lu et al. [112] developed a dual attentive GNN that can effectively predict the distribution of metro passenger flow considering the spatial and temporal influences. Further expanding on the application of attention mechanisms in GNNs, Makhdomi et al. [113] introduce a GAT method for passenger origin-destination flow prediction that leverages diverse linear and non-linear dependencies among requests originating from distinct locations, capturing both the repetition pattern and contextual data of each location. Their superior performance demonstrates the capability to capture inherent connections between regions or origin-destination pairs, and some of them [114, 112] have been deployed in the wild.
@@
Traffic resource management is to detect or schedule dedicated components in transportation systems. As an example, Wang et al. [115] and Yu et al. [116] both introduce GI-based approaches aimed at addressing critical issues in intelligent transportation systems. These frameworks capture complex dependencies within their respective applications, showcasing the versatility of GNNs in this field. However, their focuses differ significantly. Wang et al. [115] specifically design their GI model for traffic anomaly analysis in IoT-based intelligent transportation systems, aiming to identify and analyze traffic anomalies, while Yu et al. [116] focus on the operational problem of relocating and repositioning idle vehicles, a crucial aspect of traffic resource management.
V-A2 Power Grid
@@
Failure detection aims at localizing potential or yet-occurring failures within the power grid. Liao et al. [117] utilize the adjacency matrix to represent the similarity between unknown and labeled samples, and propose graph convolutional layers for identifying complex nonlinear relationships between dissolved gas and fault type. Building on the use of GCNs in power system diagnostics, Liu et al. [118] further present a technique that employs GCNs for swiftly identifying key failure points in power systems. This method simplifies and speeds up the process of detecting critical cascading failures, making it highly effective for complex power networks. Extending GI to the realm of cybersecurity in power grids, Boyaci et al. [119] introduce a deep learning model utilizing Chebyshev GCNs for detecting cyberattacks, specifically false data injection attacks, in large-scale AC power grids.
Energy harvesting prediction is for power grids with vulnerable energy input, e.g., solar energy and wind energy. Accurately predicting their future behavior is of great importance for power systems with high penetration of RESs for stable operation and economic consideration. Karimi et al. [120] focus on improving photovoltaic power forecasting by leveraging spatial and temporal coherence among power plants. They propose an STGNN to harness the relationship between power plants, observing that plants in a region experience similar environmental conditions and can thus inform each other’s power forecasts. Khodayar et al. [26] introduce a convolutional GAE for capturing continuous probability densities on graph nodes. By integrating spectral graph convolutions and variational Bayesian inference, convolutional GAE generates samples from these densities. This approach is applied to probabilistic solar irradiance prediction, using an undirected graph model of solar radiation measurement sites for future irradiance estimation. In their subsequent research, Khodayar et al. [121] further tackle the challenge of forecasting behind-the-meter load and rooftop photovoltaic generation in power systems using a spatio-temporal GAE and graph dictionary learning optimization.
V-A3 Environment Monitoring
@@
Environmental assessment and prediction aims at extracting environmental information from sensory data collected across smart cities. Chen et al. [122] introduce the group-aware GNN (GAGNN) for nationwide air quality forecasting, in order to understand latent dependencies among geographically distant cities. GAGNN constructs a hierarchical model with a city graph and city group graph, combined with a differentiable grouping network to discover and encode these complex inter-city relationships. In addition to the use of GCNs, both Gao et al. [123] and Mao et al. [124] incorporate LSTM to integrate spatio-temporal information for more accurate predictions. For example, Mao et al. [124] develop a graph convolutional temporal sliding LSTM model to predict various air pollutants. They use GCNs to model spatial dependencies and a temporal sliding LSTM strategy to capture dynamic changes over time, allowing for a broader application beyond just PM2.5 prediction.
Meteorological monitoring and forecasting leverage GI models to analyze meteorological data for understanding current weather conditions and predicting future weather patterns. Remote automatic weather stations can collect geo-distributed meteorological data for temperature prediction. The targeted sensory data are usually in a spatio-temporal form, covering atmospheric parameters such as temperature, humidity, pressure, wind speed and direction, precipitation, and other relevant variables. As an example, Lin et al. [125] present the Conditional Local Convolution Recurrent Network for spatio-temporal forecasting in weather prediction, overcoming challenges of high nonlinearity and complex spatial patterns. It introduces a GCN that captures local spatial patterns through conditional local convolutions and incorporates these into an RNN architecture to model temporal dynamics. Addressing the challenge of accurately forecasting frost events, which significantly impact agriculture, Lira et al. [126] describe a GNN with spatio-temporal attention architecture. The model, adept at handling the localized nature of frost influenced by various environmental factors, maps weather station data onto a graph structure, optimizing an adjacency matrix during training to capture complex environmental interactions. Similarly focusing on the spatial relationships among weather stations, Stanczyk et al. [127] introduce GI models for wind speed prediction, focusing on handling data from multiple weather stations.
V-A4 Public Safety